OpenStudy (anonymous):

find the difference 1/2 - 1/6 =?

OpenStudy (turingtest):

1/2-1/6=3/6-1/6=2/6=1/3

OpenStudy (anonymous):

again.... Find the common denomiator so 3/6-1/6= 3-1=2 2/8